# CALPAGAN: Calorimetry for Particles using GANs

### v1

@@ 1 Introduction
The high energy physics community, particularly with the advent of the High-Luminosity Large Hadron Collider (HL-LHC), has been increasingly engaged in the application of Deep Learning techniques, with a specific focus on Generative Adversarial Networks (GANs). Introduced by Goodfellow et al. in 2014 Goodfellow , GANs comprise two deep neural networks: the generator and the discriminator. The move towards utilizing Generative Adversarial Networks (GANs) is mainly driven by the intense computational requirements linked with the High-Luminosity Large Hadron Collider (HL-LHC). This is particularly relevant in the areas of detector modeling and the comprehensive simulation of particle interactions in the calorimeter, tasks that are recognized as some of the most demanding in terms of CPU usage.CPU . The HL-LHC’s requirements for fast and accurate simulation methods are critical, as over half of the computational resources for Large Hadron Collider experiments are dedicated to these simulations Giammanco:2014bza hep .
@@
In this study, focus shifts from the traditional full simulation approach, known for its high CPU consumption, to exploring the capabilities and limitations of a deep learning-based Generative Adversarial Network (GAN). This is aimed at achieving both high precision and rapid calorimeter simulations. The goal is to develop an innovative fast simulation technique that could significantly reduce both computation time and disk space requirements in LHC and future physics research. paganini
@@
The structure of this paper is organized as follows: Section 1.1 offers a comprehensive description of the simulation and reconstruction processes for the input images (condition) as well as the target images. Section 2 details the architecture of both the generator and discriminator networks. Subsequently, Section 3 presents a performance evaluation of the generated outcomes by comparing the distributions of high-level jet properties.
1.1 Input Preparation
@@
Monte Carlo datasets were generated utilizing CMSSW, the software framework designed for CMS. This extensive procedure encompassed several stages: Generation, Simulation, Digitization, and Reconstruction, collectively known as GEN-SIM-DIGI-RECO chain. The event simulation was carried out using both Delphes and Geant4.
@@
In this research, GEANT4 and Madgraph4 combined with Pythia8 and Delphes were independently used to generate 10,000 events each of W+Jet (with up to 3 jets) and dijet. To ensure the generation of identical events, both simulations utilized the same Pythia8 output, which was stored in “HepMC” format files. The dataset was divided into two equal parts: 5,000 events for training and 5,000 for testing. Subsequently, two-dimensional calorimeter images were produced by summing the hadronic and electromagnetic energy in the calorimeter towers. The calorimeter structure of the CMS Detector allows the creation of a $72\times 72$ image in eta phi space. Moreover, since the pixel values in these images represent energy rather than color (as shown in Fig. 1), they are single-channel. This characteristic means that the standard pix2pixpix2pix GAN architecture cannot be directly applied, thus necessitating some modifications to the network structure.
@@
Figure 1: A calorimetric energy image in $\eta$- $\phi$ space (respectively horizontal and vertical axis) of an event generated using GEANT4.
@@
Calorimeter images produced by Delphes serve as inputs for the generator. This generator creates random distributions using the calorimetric jet data from Delphes. On the other hand, the discriminator will decide how realistic a distribution the generator has produced thanks to simulations performed by Geant4 as illustrated in Fig. 2.
Figure 2: The GAN architecture used for training CALPAGAN.
@@ 2.1 The CALPAGAN
The model employed in this study, as shown in Fig. 3, known as cGAN (conditional GAN), is designed to transform the input image, which also serves as the condition, into the target image.
Figure 3: The applied model.
In the Pix2pix framework, the generator’s architecture is based on U-Net. As illustrated in Fig. 4, U-Net structure first compresses the input image down to a bottleneck size, and then, following the bottleneck, it expands the image back to the required output size.
Figure 4: Architecture of the generative network (U-Net).
@@
The images from Delphes were normalized by dividing each pixel value by the highest-energy pixel value found within the Delphes images themselves. Similarly, the images generated by Geant4 were normalized by dividing each pixel value by the highest-energy pixel value in the Geant4 images. Previously, when pixel values were normalized to the range of [-1, 1], pixels with zero energy reduced the loss function to -1. This led to the generator producing blank images after a few epochs.
The loss function for the CALPAGAN is composed of two distinct components. The first is the conventional loss function typically employed in cGANs. The second component is the L1 loss function, also known as the mean absolute error (MAE) loss function. This is calculated as the average of the absolute differences and is particularly effective in reducing blurriness in the images (as shown in Fig. 5). By integrating these two loss functions, we formulated the following loss function, which was, of course, aimed to be minimized during the training process.
Figure 5: Loss function of the CALPAGAN
The jets were detected and identified using the FastJet program fastjet , which utilizes the energy deposits in the calorimeter towers and the outcomes from the Geant4 simulation. Subsequently, a trimming process, as described by Krohn et al. in 2010 jettrimming , was employed to remove energy contributions from pile-up events.
Figure 6: The comparison of event images obtained from the test set by CALPAGAN and their corresponding event images from Delphes and Geant4. The image at the top is a randomly selected example from 5,000 test events. The one below shows the jets obtained from these images in $\eta$ (horizontal) and $\Phi$ (vertical) space.
@@ 3 Results and Conclusion
Inspired by the pix2pix model, we developed the CALPAGAN model with the aim of creating calorimeter images similar to those produced by GEANT4, using Delphes images as a basis. CALPAGAN has been effective in producing images very close to those from GEANT4, thereby reducing computational costs. However, during this research, various challenges were encountered, the most significant of which was the natural sparsity observed in the data.
Data sparsity emerged as a significant limiting factor for the success of CALPAGAN. The presence of numerous zero-valued pixels in the calorimeter images adversely affected weight calculations. Despite testing with various activation functions such as ReLU, sigmoid, and softmax, as well as experimenting different filter sizes and strides within the generator network, these attempts failed to yield positive results.
Pix2pix is successful in creating synthetic images that closely resemble real images. For example, when tasked with creating artificial tree images, a GAN based on pix2pix outperforms a GAN designed to mimic images created by particles dispersing energy at specific points in the calorimeter. Nonetheless, visually identifying the similarities between synthetic calorimeter images and their real counterparts remains a complex task. One widely recognized metric for evaluating GANs is the Fréchet Inception Distance (FID)FID . Consequently, FID was employed to assess the similarity between the generated and real images.
@@
In Fig.9, the distributions of 1-subjettiness $\tau_{1}$ and 2-subjettiness $\tau_{2}$ obtained from GEANT4, Delphes, and CALPAGAN are shown for the first and second jets. n-subjettiness, designed to identify boosted objects that decay hadronically, such as electroweak bosons and top quarks subjettiness , is particularly suitable for the Delphes-GEANT4-CALPAGAN comparison due to its variable nature that captures the internal structure of jets.
@@
Evaluating the speed enhancement of CALPAGAN in comparison to GEANT4 was one of our primary objectives. Although given similar acceleration rates in similar studies, to make this comparison directly, the proposed model should provide all the outputs offered by the GEANT4 simulation. Therefore, instead of comparing the inference time required for an event of the model we obtained, it was obtained for different batch size values for CPU, single GPU, and dual GPU. As seen from the graph in Fig.Fig.15, GPU usage becomes more advantageous than CPU as the batch size increases. Of course, it is also important to note that the upper limit of batch size is RAM size (some may use swap memory with GPU, but in practice, this will increase inference time even more). According to the results obtained in this study, the use of two GPUs makes the fastest inference per event when the batch size value is 512.

### v2

@@ 1 Introduction
The high energy physics community, particularly with the advent of the High-Luminosity Large Hadron Collider (HL-LHC), has been increasingly engaged in the application of deep learning techniques, with a specific focus on Generative Adversarial Networks (GANs) musella . Introduced by Goodfellow et al. in 2014 Goodfellow , GANs comprise two deep neural networks: the generator and the discriminator. The move towards utilizing Generative Adversarial Networks (GANs) is mainly driven by the intense computational requirements linked with the High-Luminosity Large Hadron Collider (HL-LHC). This is particularly relevant in the areas of detector modeling and the comprehensive simulation of particle interactions in the calorimeter, tasks that are recognized as some of the most demanding in terms of CPU usage.CPU . The HL-LHC’s requirements for fast and accurate simulation methods are critical, as over half of the computational resources for Large Hadron Collider experiments are dedicated to these simulations Giammanco:2014bza hep .
@@
In this study, focus shifts from the traditional full simulation approach, known for its high CPU consumption, to exploring the capabilities and limitations of a deep learning-based GAN. This is aimed at achieving both high precision and rapid calorimeter simulations. The goal is to develop an innovative fast simulation technique that could significantly reduce both computation time and disk space requirements in LHC and future physics research. paganini
@@
The structure of this paper is organized as follows: Section 2 offers a comprehensive description of the simulation and reconstruction processes for the input images (condition) as well as the target images. Section 3 details the architecture of both the generator and discriminator networks. Subsequently, Section 4 presents a performance evaluation of the generated outcomes by comparing the distributions of high-level jet properties.
2 Calorimeter data simulation and input data pre-processing
@@
Monte Carlo datasets were generated for CMS Detector by utilizing CMSSWcmssw , the software framework designed for CMS. CALPAGAN needs a Delphes image as input along with a noise, and the target image is the image obtained from GEANT4 simulation which is only possible using CMSSW. This extensive procedure encompassed several stages: Generation, Simulation, Digitization, and Reconstruction, collectively known as GEN-SIM-DIGI-RECO chain. The event simulation was carried out using both Delphes and Geant4.
@@
In this research, Geant4 and Madgraph4 madgraph4 combined with Pythia8 and Delphes delphes were independently used to generate 10,000 events each of W+Jet (with up to 3 jets) and dijet. To ensure the generation of identical events, both simulations utilized the same Pythia8 output, which was stored in “HepMC” format files. The dataset was divided into two equal parts: 5,000 events for training and 5,000 for testing. Subsequently, two-dimensional calorimeter images were produced by summing the hadronic and electromagnetic energy in the calorimeter towers. The calorimeter structure of the CMS Detector allows the creation of a $72\times 72$ image in $\eta-\phi$ space. Moreover, since the pixel values in these images represent energy rather than color (as shown in Fig. 1), they are single-channel. This characteristic means that the standard Pix2pixpix2pix GAN architecture cannot be directly applied, thus necessitating some modifications to the network structure. Pix2pix is successful in creating synthetic images that closely resemble real images. For example, when tasked with creating artificial tree images, a GAN based on Pix2pix outperforms a GAN designed to mimic images created by particles dispersing energy at specific points in the calorimeter.
@@
Figure 1: This image displays the calorimetric energy in GeV of an event created with Geant4, presented on a 72 x 72 grid (5184 pixels total) within the $\eta$-$\phi$ space. Here, the horizontal axis represents the $\eta$ indices and the vertical axis represents the $\phi$ indices respectively.
As seen in Fig. 2, z, the random noise, has been added to calorimeter images produced by Delphes, x, then served as input for the generator G. The generated data G(x) is then added pixel by pixel to x, producing a combined input G(x)+x for the discriminator D. The discriminator evaluates both this combined input and the real samples y, learning to distinguish between real and counterfeit data through iterative training.
@@
The generator of CALPAGAN creates random distributions using the calorimetric jet data from Delphes. Finally, the discriminator network (D) decides how realistic the distribution produced by the generator has produced thanks to simulations performed by Geant4 as illustrated in Fig. 2.
Figure 2: Training architecture of CALPAGAN: x represents the input images, which are calorimeter images produced by Delphes. z is random noise. G(x) denotes the generated data, while y signifies the real images which is calorimeter images produced by Geant4. D represents the discriminator.
Actually, this is a game played by the generator and the discriminator, known as a zero-sum game Nash in the literature.The generator network creates data and engages in a zero-sum game with a discriminator network, aiming to reach a Nash equilibrium Nash based on the feedback from the discriminator. At equilibrium, the discriminator must randomly guess with a 50 chance whether the generator’s data is real or fake.
@@ 3.1 The CALPAGAN
The model employed in this study, as shown in Fig., known as cGAN (conditional GAN)cGAN , is designed to transform the input image, which also serves as the condition, into the target image.
In the Pix2pix framework, the generator’s architecture is based on U-NetOlaf . U-Net is an advanced deep learning framework that utilizes an encoder-decoder structure with convolutional networks to efficiently process and analyze data. Its unique feature, the incorporation of skip connections, ensures the preservation of important spatial information. As illustrated in Fig. 3, U-Net structure first compresses the input image down to a bottleneck size, and then, following the bottleneck, it expands the image back to the required output size.
Figure 3: The U-Net architecture of the generator of CALPAGAN.
@@
Additionally, the Delphes calorimeter images used in the training process were normalized by dividing by the highest energy pixel value of each Delphes image, and similarly, the Geant4 images were normalized by the highest energy pixel value of each Geant4 image. This normalization moved pixel values to the range [0, 1], which also partially addressed the sparsity problem by eliminating the impact of zero-energy pixels on the loss function (as opposed to when pixel values were previously normalized to the range [-1, 1], where zero-energy pixels would contribute a value of -1 to the loss function, leading to the generator producing blank images after a few epochs.).
The loss function for the CALPAGAN is composed of two distinct components. The first is the conventional loss function typically employed in cGANs. The second component is the L1 loss function, also known as the mean absolute error (MAE) loss function. This is calculated as the average of the absolute differences and is particularly effective in reducing blurriness in the images as shown in below. By integrating these two loss functions, we formulated the following loss function 1 , which was, of course, aimed to be minimized during the training process. In the formula, x, y, and z represent the input image, output image, and random noise, respectively.
$\mathcal{L}_{cGAN}(G,D)$ $=\mathbb{E}_{x,y}[\log D(x,y)]+\mathbb{E}_{x,z}[\log(1-D(x,G(x,z)))]$ (1)
$\mathcal{L}_{L1}(G)$ $=\mathbb{E}_{x,y,z}[\|y-G(x,z)\|_{1}]$
$G^{*}$ $=\arg\min_{G}\max_{D}\mathcal{L}_{cGAN}(G,D)+\lambda\mathcal{L}_{L1}(G)$
The jets were detected and identified using the FastJet program fastjet , which utilizes the electromagnetic and hadronic energy depositions in the calorimeter(ECAL+HCAL) towers and the outcomes from the Geant4 simulation. Subsequently, a trimming process, as described by Krohn et al. in 2010 jettrimming , was employed to remove energy contributions from pile-up events.
Figure 4: Comparison of event images obtained from the test set with Delphes and Geant4 event images (top: a randomly selected example from 5000 test events). The bottom image shows the images of jets obtained from these event images in the $\eta$ (horizontal) and $\Phi$ (vertical) space.
@@ 4 Results and Conclusion
We first tried to apply Pix2pix architecture as it is. However, it didn’t perform so well. So, we have tweaked the architecture to get better results. Inspired by the Pix2pix, CALPAGAN has been developed with the aim of creating calorimeter images similar to those produced by Geant4, using Delphes images as a condition. CALPAGAN has been effective in producing images very close to those from Geant4, thereby reducing computational costs. However, during this research, various challenges were encountered and the most significant of which was the natural sparsity observed in the data.
Data sparsity limits the success of CALPAGAN. The presence of numerous zero-valued pixels in the calorimeter images adversely affected weight calculations. Despite testing with various activation functions such as ReLU, sigmoid, and softmax, as well as experimenting different filter sizes and strides within the generator network, these attempts failed to yield positive results.
Nonetheless, visually identifying the similarities between synthetic calorimeter images and their real counterparts remains a complex task. The widely recognized metric for evaluating GANs is the Fréchet Inception Distance (FID)FID . Consequently, FID was employed to assess the similarity between the generated and real images.
@@
Additonally, the Fig. 6 shows the change in the loss function values obtained from the test sample in the GAN training conducted with jet images as the number of epochs progresses.
The loss value of the Generator in a CALPAGAN remains continuously high, and the image generation performance is relatively low, it typically suggests that the Generator is struggling to produce realistic images that can fool the Discriminator.
In our CALPAGAN model, inspired by the Pix2pix architecture, we observed that the discriminator learns rapidly, outpacing the generator and resulting in unstable training dynamics. This quick learning provides overly strong gradient feedback, making it difficult for the generator to catch up. Additionally, the architectural imbalance between the simpler discriminator and the more complex generator makes the problem worse, leading to poor performance. This issue can also stem from the natural sparsity of the data. Consequently, designing a GAN to mimic images of energy dispersion by particles at specific locations within a calorimeter becomes challenging.
The red line represents the loss of the discriminator. It is much steadier and remains low, which implies that the discriminator is performing its task with relative ease. The low and steady loss suggests that the discriminator quickly becomes good at distinguishing real images from the fake ones generated by the generator.
Therefore the generator might need a more complex architecture to improve its performance. here might be a need for techniques to stabilize the training, such as modified loss functions, regularization methods, or different training strategies which we have already tried some of them.
In Fig.7, the distributions of 1-subjettiness $\tau_{1}$ and 2-subjettiness $\tau_{2}$ obtained from Geant4, Delphes, and CALPAGAN are shown for the first and second jets.’N-subjettiness– designed to identify boosted hadronically-decaying objects like electroweak bosons and top quarks. Combined with a jet invariant mass cut, N-subjettiness is an effective discriminating variable for tagging boosted objects and rejecting the background of QCD jets with large invariant mass’ subjettiness , is particularly suitable for the Delphes-Geant4-CALPAGAN comparison due to its variable nature that captures the internal structure of jets.
@@
The linear radial moment or girth is a specific case where a general radial moment is defined as f(r)=r girth . It is defined based on the components that constitute a jet as follows:
$g=\sum_{i\in jet}\frac{(p^{i}_{T})}{p^{jet}_{T}}{r_{i}}$
@@
Evaluating the speed enhancement of CALPAGAN in comparison to Geant4 was one of our primary objectives. Although given similar acceleration rates in similar studies, to make this comparison directly, the proposed model should provide all the outputs offered by the Geant4 simulation. Therefore, instead of comparing the inference time required for an event of the model , it was obtained for different batch size values for CPU, single GPU, and dual GPU. For the dual GPU setup, we used a model parallelism approach implemented in PyTorch.
As it can be seen in Fig.13, GPU usage becomes more advantageous than CPU as the batch size increases. Of course, it is also important to note that the upper limit of batch size is RAM size (some may use swap memory with GPU, but in practice, this will increase inference time even more).
Moreover, as seen in 13, the dual GPU usage becomes advantageous for batch sizes 128 and more. This effect is most probably due to the data transfer latency. When using multiple GPUs, data needs to be transferred between them. For smaller batch sizes, the time taken to transfer data can overshadow the benefits of parallel computation. According to the results, the use of dual GPU makes the fastest inference per event when the batch size value is 512.